A uniform, spherical cloud of interstellar gas has mass 2.2×1030 kg and radius 9.0×1012 m , and is rotating with period 1.5×106 years.

If the cloud collapses to form a star 7.4×108 m in radius, what will be the star's rotation period?
Answer #1

From the conservation of angular momentum,

We Write

I1W1 = I2*W2

I sphere = 2/5 *M*R^2

So,

Wf = (Ri/Rf)^2 *Wi

Wf = (7.4*10^8/ 9*10^12)^2 * 1.5*10^6

Wf = 0.01014 Years

Star's Rotation Period = 0.01014 Years = 3.701 days